A solution is 20% by weight in H2SO4 and has a density of 1.14g/mL. Determines: (a) Molarity of solution b) Molarity of solution
1
Expert's answer
2020-09-23T04:44:13-0400

20% by weight means 20 g of H2SO4 present in 100g solution.

moles of H2SO4 = "\frac{20}{98} moles"

Molarity = no. of moles of solute/vol. of solution

Molarity = "\frac{\frac{20}{98} }{\frac{100}{1000}}"

Molarity = 2.04 M



density of solution is given = 1.14g/mL

so mass of solution = 100*1.14 = 114 g

mass of solvent = 114 - 20 g = 94 g


Molality = no. of moles of solute / mass of solvent

Molality = "\frac{\frac{20}{98} }{\frac{94}{1000}}"

Molality = 2.17 molal